how to find the critical value for an alpha level .001 for a 2-tailed test?
Answer #1

Solution :

Given that,

This is the two tailed test .

\alpha = 0.001

\alpha / 2 = 0.001 / 2 = 0.0005

P(Z > z) = 0.001

1 - P(Z < z) = 0.0005

P(Z < z) = 1 - 0.0005

P(Z < z) = 0.9995

P(Z < 3.29) = 0.9995

Critical value = 3.29